need help with these problems. im stuck. please solve in detail and provide your answer. i will rate fast and high
#1

1.-(50) Find: Vas Im I and the region of Ro RG 15V NW 414 geration 60 mA/ 4:2V PG IMA R = 200.2
#2

2-(50) Find Ib, Ic , Yce and the region of operation 200.12 IM 2 B= 20 200
Answer #1

1 Given: Rolma] Ro=loor #14 is v voi Transistor (MOSFT) operates in saturation region. since, yo=v6 becau IG=0. So, IG RG =0ID kn [vois vt]² kn = 60mA/v2 Vt = 2V .ID = GO MA/z [[ugs-2] ² But VGS 2 on VOT*. I. -51 ID = 30mA [ 133 - In 200]? 12 = ID- ID2 40000 5233:33 Io +1690 After solving Io = 72.78 mA and 58.84 mA To check the valid root =ON Io 200 root 1 sl=an 14.556VGS = VDs = is - Ig(200) VGS 15 - 200 (58.04 m47) VGS 3.92v. T m [2 kn ID = /2 * somn *58.04MA gm = 16 964.8 mo gm= .83.45 mv3) 200 22 3V. IMA I M p=200 UBE - For IB KVL to the input Loop. V + (FB+7c) 200+ IB IM TUBE TO Assume UBE = 0.7 - IB ( 1 + B)IB = 1.64 MA ☺ Ic Ic= BIB Ic= 200 (1.64 MA) Ic= 3281MA I c = 328. led calculate UBE and VCEVCE VCE= Vc -VE VC 3-(Ict IB) 200 3 - (1.64+328-1uA 200 3 - 65.82 mv = 3 ooo6582v. VCE 239341region ca operation since, VBE LUCE the transistor operates in Active region.